# Invert4TVG: A Temporal Video Grounding Framework with Inversion Tasks for Enhanced Action Understanding

### v1

@@ Introduction
IoU Reward. As stated above, TVG aims at estimating the time interval in the video that is associated with the content of a given textual query. We use the Intersection over Union (IoU) (Yuan et al. 2021) between the time interval predicted by the model and the ground-truth interval as the reward function. This reward function effectively describes the accuracy of the time interval predicted by the model. Given predicted $t_{s},t_{e}]$ and ground-truth $t^{\prime}_{s},t^{\prime}_{e}]$ segments, the IoU reward can be calculated as follows:
@@
Format Reward. Recently, Time-R1 (Xu 2025a) and VideoChat-R1 (Li et al. 2025) employ a Format reward making the model explicitly output its thinking process before making predictions. Following them, we introduce a template-based reasoning reward that incentivizes the model to generate intermediate reasoning steps prior to providing answers. The format is as following: <think>***</think> <answer> $t_{s}$ to $t_{e}$ </answer>. The reward is formulated as:
@@
Invert4TVG Reinforcement Learning Framework
While a joint training approach that processes all TVG and Invert-TVG tasks simultaneously might seem straightforward, this method suffers from several critical limitations: (1) Memory inefficiency: maintaining separate computation graphs for multiple tasks drastically increases GPU memory consumption; (2) Optimization conflict: gradient updates from different tasks may interfere with each other, especially when their loss landscapes are not aligned; (3) Training instability: the varying convergence rates of different tasks make it challenging to balance their contributions; (4) Task bias: the model may prioritize easier tasks while neglecting others. These drawbacks motivate our proposed alternating training paradigm.
We (see Figure 3) implement a probabilistic sampling strategy where each training iteration has an high probability (80% in this paper) of executing the primary TVG task (using IoU and format rewards) and a low probability of performing an Invert-TVG task. When selecting Invert-TVG, we uniformly sample among VC, AR and VD. This design ensures the model maintains its core action understanding capabilities while primarily focusing on temporal localization. The asymmetric probability distribution prevents the auxiliary tasks from overwhelming the main objective while still providing regular semantic reinforcement. Formally, the reward for training the TVG task is:
@@
As stated above, $p$ is 0.8 in this paper.
Experiments
@@
Table 1: Performance of temporal video grounding on Charades-STA. The methods marked with * indicate the use of additional VTG datasets, while FT with ✓indicates training on the Charades-STA training set. We compare our method against existing 3B, 7B open-source LVLM, as well as state-of-the-art VLP models. We highlight our results and the best-performing baselines using bold and underlining for clear comparison.
@@
Figure 4: Performance of temporal video grounding on ActivityNet and Qvhighlight (RL methods). All the models are Zero-Shot on the test dataset.
Experimental Setup
Benchmarks. We test our model on three temporal video grounding datasets: (1) Charades-STA (Sigurdsson et al. 2016) contains 6,672 long videos capturing indoor human activities. The official split for the TVG task includes 12,408 clip-query pairs for training and 3,720 for testing. (2) ActivityNet (Heilbron et al. 2015) comprises 20K long videos with an average of 3.65 clip-query pairs per video. We use the standard dataset splits with 37,421 training, 17,505 validation, and 17,031 test samples. (3) We further evaluate on QvHighlight (Lei, Berg, and Bansal 2021), a high-resolution set of 10,460 long YouTube videos paired with 48k manually annotated clip-queries. To match Charades-STA and ActivityNet formats, multi-segment localizations are split into single-segment tasks, forming a balanced benchmark for fine-grained temporal grounding.
Implementation Details We implement our LVLM using the Qwen2.5-VL model (Bai et al. 2025) as the backbone. To balance efficiency and memory consumption, we sample video frames at 2 FPS and resize them, resulting in approximately 2.8 million pixels per video (e.g., a 50-second video yields 100 frames of size $96\times 96\times 3$). Our implementation utilizes SpaCy’s en_core_web_sm-3.8.0 model (12MB) to extract verbs from sentences. For optimization, we employ the AdamW optimizer (Loshchilov and Hutter 2019) with the following parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$, a weight decay of 0.0, and a learning rate of $5\times 10^{-5}$. The training time per epoch is approximately 80 hours. To ensure reproducibility, our code, configuration files, and execution scripts are available in the supplementary materials.
Comparison with State-of-the-Art
@@
We compare Invert4TVG with state-of-the-art TVG methods, including both traditional video-language pre-training models (VLP), recent large video-language models fine-tuned via SFT and other RL-based approaches.
@@
Invert4TVG surpasses other models on the Charades-STA dataset. As shown in Table 1, Invert4TVG surpasses not only VLP-based and SFT-based models but also outperforms RL-based approaches under identical conditions. For example, on Charades-STA, the 7B variant of Invert4TVG achieves an R1@0.7 of 51.4, exceeding TimeSuite (43.0), SnAG (46.2), and Time-R1 (50.1). The improvements are more pronounced for the 3B variant. Across R1@0.3, R1@0.5, and R1@0.7, Invert4TVG’s 3B model consistently outperforms the 3B version of Time-R1.
Invert4TVG surpasses other RL-based models in zero-shot settings. As shown in Figure 4, in zero-shot settings, Invert4TVG’s 3B and 7B variants also outperform Time-R1 on R1@0.3, R1@0.5, and R1@0.7 over ActivityNet. On QvHighlight, where we compare single-segment predictions, Invert4TVG consistently outperforms Time-R1 across R1@0.3, R1@0.5, and R1@0.7, demonstrating its superior accuracy in complex scenarios.
Ablation Study
@@
We conduct a detailed ablation on the Invert4TVG-3B model to investigate the contribution of various design strategies.
@@
Using a single Invert-TVG task versus employing them in combination. As shown in Table 2, using VC, AR, or VD alone instead of jointly yields lower performance. Only-VD, which emphasizes contextual understanding, peaks at R1@0.3 but falls short on precise localization. Only-AR, focused on immediate actions, reaches the highest R1@0.7 of 43.8. Only-VC outputs are less random than Only-VD yet less specific than Only-AR, achieving the best R1@0.5 (68.0). The mixed-task Invert4TVG surpasses all three individual tasks across all three metrics, demonstrating that joint training outperforms separate use.
Exploiting different probabilities for the TVG and Invert-TVG tasks. As shown in Figure 5, Varying the task probability markedly alters training outcomes. At 20% Invert-TVG task probability, the model performs best, raising R1@0.7 from 36.9 (no Invert task) to 44.0. As the Invert-TVG task probability grows, the model increasingly emphasizes action recognition while neglecting temporal grounding. Between 60% and 80% Invert-TVG, temporal video grounding performance steadily declines, falling below the pure TVG baseline. When Invert-TVG probability reaches 100 %, the model performs only the Invert-TVG task and yields the worst results.
@@
Table 2: Ablation study using only VC (Verb Completion), AR (action recognition), VD (video description), and their mixed usage.
Figure 5: The R1 accuracy curves. Blue, orange, and green show how the three R1 metrics evolve as the Invert-TVG task probability gradually increases.
Binary Invert-TVG rewards vs. cosine similarity-based rewards. As shown in Table 3, we observe that employing a simple binary Invert-TVG reward (0 or 1) during training yields superior outcomes compared to more intricate reward mechanisms. At identical training steps, the employed Invert-TVG Reward outperforms the cosine similarity reward across all three evaluation metrics (R1@0.3, R1@0.5, R1@0.7). This advantage stems from the controllability and stability of the binary reward design, whereas cosine similarity introduces higher variance and optimization instability. For example, “run” and “eat” yield a cosine similarity of 0.2 despite their weak semantic link. Therefore, binary Invert-TVG reward is a better choice.
@@
Discussion: Why Invert-TVG Tasks? Is there any alternative?
While our framework proposes three Invert-TVG tasks (VC, AR, VD), other candidates could theoretically enrich video-text understanding. For instance, contrastive learning between video segments and queries could enforce global alignment, or masked frame modeling might enhance temporal reasoning. However, these alternatives lack the task-specific synergy and semantic granularity that make VC, AR, and VD indispensable for TVG. In our method, VC, AR, and VD are directly derived from the TVG task itself, each engineered to isolate and strengthen a critical dimension of action understanding that directly impacts temporal localization. In essence, VC, AR, and VD are not generic self-supervised tasks but purpose-built derivatives of the TVG inverse problem. They distill the essence of action understanding from the TVG task itself, ensuring immediate and targeted improvements in localization accuracy. While contrastive learning or masked modeling may enhance broad semantic capabilities, their indirect impact on TVG-specific action-query alignment lacks the precision and immediacy of our task-specific design.
Conclusion

### v2

@@ 2.3 IoU and Format Reward Functions
IoU Reward. As stated above, TVG aims at estimating the time interval in the video that is associated with the content of a given textual query. We use the Intersection over Union (IoU) (Yuan et al., 2021) between the time interval predicted by the model and the ground-truth interval as the reward function. This reward function effectively describes the accuracy of the time interval predicted by the model. Given predicted $t_{s},t_{e}]$ and ground-truth $t^{\prime}_{s},t^{\prime}_{e}]$ segments, the IoU reward can be calculated as follows:
@@
Format Reward. Recently, Time-R1 (Wang et al., 2025a) and VideoChat-R1 (Li et al., 2025) employ a Format reward making the model explicitly output its thinking process before making predictions. Following them, we introduce a template-based reasoning reward that incentivizes the model to generate intermediate reasoning steps prior to providing answers. The format is as following: <think>***</think> <answer> $t_{s}$ to $t_{e}$ </answer>. The reward is formulated as:
@@
2.4 Invert4TVG Reinforcement Learning Framework
While a joint training approach that processes all TVG and Invert-TVG tasks simultaneously might seem straightforward, this method suffers from several critical limitations: (1) Memory inefficiency: maintaining separate computation graphs for multiple tasks drastically increases GPU memory consumption; (2) Optimization conflict: gradient updates from different tasks may interfere with each other, especially when their loss landscapes are not aligned; (3) Training instability: the varying convergence rates of different tasks make it challenging to balance their contributions; (4) Task bias: the model may prioritize easier tasks while neglecting others. These drawbacks motivate us to adopt the training paradigm illustrated in Figure 3.
We implement a probabilistic sampling strategy where each training iteration has a high probability (80% in default) of executing the primary TVG task (using IoU and format rewards) and a low probability of performing an Invert-TVG task. When selecting Invert-TVG, we uniformly sample among VC, AR and VD. This design ensures the model maintains its core action understanding capabilities while primarily focusing on temporal grounding. The asymmetric probability distribution prevents the auxiliary tasks from overwhelming the main objective while still providing regular semantic reinforcement. Formally, the reward for training the TVG task is:
@@ 2.4 Invert4TVG Reinforcement Learning Framework
As mentioned above, $p=0.8$ is an empirically determined parameter derived from experiments.
3 Experiments
@@
3.1 Experimental Setup
Benchmarks. We test our model on three temporal video grounding datasets: (1) Charades-STA (Sigurdsson et al., 2016) contains 6,672 long videos capturing indoor human activities. The official split for the TVG task includes 12,408 clip-query pairs for training and 3,720 for testing. (2) ActivityNet (Heilbron et al., 2015) comprises 20K long videos with an average of 3.65 clip-query pairs per video. We use the standard dataset splits with 37,421 training, 17,505 validation, and 17,031 test samples. (3) We further evaluate on QvHighlight (Lei et al., 2021), a high-resolution set of 10,460 long YouTube videos paired with 48k manually annotated clip-queries. To match Charades-STA and ActivityNet formats, multi-segment localizations are split into single-segment tasks, forming a balanced benchmark for fine-grained temporal grounding.
Implementation Details. We implement our LVLM using the Qwen2.5-VL model (Bai et al., 2025) as the backbone. To balance efficiency and memory consumption, we sample video frames at 2 FPS and resize them, resulting in approximately 2.8 million pixels per video (e.g., a 50-second video yields 100 frames of size $96\times 96\times 3$). Our implementation utilizes SpaCy’s en_core_web_sm-3.8.0 model (12MB) to extract verbs from sentences and transform them across different tenses. For optimization, we employ the AdamW optimizer (Loshchilov and Hutter, 2019) with the following parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$, a weight decay of 0.0, and a learning rate of $5\times 10^{-5}$. The training time per epoch is approximately 80 hours. To ensure reproducibility, our code, configuration files, and execution scripts are available in the supplementary materials.
Table 1: Performance of temporal video grounding on Charades-STA. Methods labeled FT with a ✓were fine-tuned on the Charades-STA training set. Methods marked with * were first pre-trained on extra TVG datasets1 and then fine-tuned on the Charades-STA training set, while those without * are only trained on Charades-STA. We compare our method against existing 3B, 7B open-source LVLM. We highlight our results and the best-performing baselines using bold and underlining for clear comparison.
@@ 3.1 Experimental Setup
1 YT-Temporal, DiDeMo, QuerYD, InternVid, HowTo100M datasets, our method is not pretrained on those datasets.
@@
Figure 4: Performance of temporal video grounding on ActivityNet and QvHighlight. We compare our method with Time-R1 (the best-performing among previous methods). All the models are zero-shot tested.
3.2 Comparison with State-of-the-Art Approaches
@@
We compare Invert4TVG with state-of-the-art TVG methods, including both traditional video-language pre-training models (VLP), recent large video-language models fine-tuned via SFT and RL-based approaches.
@@
Comparisons on the Charades-STA dataset with fine-tuning. As shown in Table 1, Invert4TVG surpasses not only VLP-based and SFT-based models but also outperforms RL-based approaches under identical conditions. For example, on Charades-STA, the 7B variant of Invert4TVG achieves an R1@0.7 of 51.4, exceeding TimeSuite (43.0), SnAG (46.2), and Time-R1 (50.1). The improvements are more pronounced for the 3B variant. Across R1@0.3, R1@0.5, and R1@0.7, Invert4TVG’s 3B model outperforms the 3B version of Time-R1.
Comparisons on the ActivityNet and QvHighlight datasets in zero-shot settings. As shown in Figure 4, in zero-shot settings, Invert4TVG’s 3B and 7B variants outperform Time-R1 on R1@0.3, R1@0.5, and R1@0.7 over ActivityNet. On QvHighlight, where we compare single-segment predictions, Invert4TVG consistently outperforms Time-R1 across R1@0.3, R1@0.5, and R1@0.7. ActivityNet contains only 200 action categories, whereas QvHighlight covers a significantly larger and more diverse set, with far more complex scene–action correlations. This disparity underscores the superiority of our method in understanding intricate actions.
3.3 Ablation Study
@@
We conduct a detailed ablation on the Invert4TVG-3B model to investigate the contribution of the design strategies.
@@
Using different combinations of Invert-TVG tasks versus employing them in combination. As shown in Table 2, using VC, AR, or VD alone instead of jointly yields lower performance. Only-VD, which emphasizes contextual understanding, peaks at R1@0.3 but falls short on precise localization. Only-AR, focused on immediate actions, reaches the highest R1@0.7 of 43.8. Only-VC outputs are less random than Only-VD yet less specific than Only-AR, achieving the best R1@0.5 (68.0). The mixed-task Invert4TVG surpasses all three individual tasks across all three metrics, demonstrating that joint training outperforms separate use.
Table 2: Ablation study using only TVG (Time-R1), VC (verb Completion), AR (action recognition), VD (video description), and their mixed usage.
@@
The combination of VC and AR improves R1@0.7 to 43.8, outperforming either task alone (VC: 42.0; AR: 43.8), indicating complementary benefits between verb completion and action recognition. The VC+VD pair achieves the highest R1@0.3 (80.0) among two-task setups, suggesting that video description aids verb-focused localization. Invert4TVG (integrating all three auxiliary tasks) achieves the best overall results (R1@0.7: 44.0), demonstrating that multi-task synergy is maximized when all components are jointly optimized.
Figure 5: The R1 accuracy curves. Blue, orange, and green show how the three R1 metrics evolve as the Invert-TVG task probability (1p)(1-p) gradually increases.
Exploiting different probabilities for the TVG and Invert-TVG tasks. As shown in Figure 5, Varying the task probability markedly alters training outcomes. At 20% Invert-TVG task probability, the model performs best, raising R1@0.7 from 36.9 (no Invert task) to 44.0. As the Invert-TVG task probability grows, the model increasingly emphasizes action recognition while neglecting temporal grounding. Between 60% and 80% Invert-TVG, temporal video grounding performance steadily declines, falling below the pure TVG baseline. When Invert-TVG probability reaches 100 %, the model performs only the Invert-TVG task and yields the worst results. According to the experiments, we choose to set $p=0.8$ in Eq. 13.
@@
Binary Invert-TVG rewards vs. cosine similarity-based rewards. As shown in Table 3, we observe that employing a simple binary Invert-TVG reward (0 or 1) during training yields superior outcomes compared to more intricate reward mechanisms. When training for the same two epochs, the employed Invert-TVG Reward outperforms the cosine similarity reward across all three evaluation metrics (R1@0.3, R1@0.5, R1@0.7). This advantage stems from the controllability and stability of the binary reward design, whereas cosine similarity introduces higher variance and optimization instability. For example, in our implementation, “run” and “eat” yield a cosine similarity of 0.2 despite their weak semantic link. Therefore, binary Invert-TVG reward is a better choice.
4 Conclusion
@@
Acknowledgments
This research was financially supported by the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) under Grant No. GML-KF-24-17, Guangdong Provincial Basic and Applied Basic Research Fund under Grant No. 2025A1515011884, and Fundamental Research Funds for the Central Universities under Grant No. 2024ZYGXZR021.
ETHICS STATEMENT
This study exclusively utilizes publicly available open-source models and datasets; no proprietary or sensitive information is involved, and all data are free of personally identifiable content. We have strictly followed the corresponding licenses and usage guidelines. Although the present work poses no apparent ethical risks, we caution that—like many machine learning models—its outputs could be misapplied in unforeseen contexts. We therefore advocate responsible use and encourage ongoing efforts to identify and mitigate potential biases inherent in open-source datasets.
REPRODUCIBILITY STATEMENT
We are committed to ensuring the reproducibility of our work, The models, training datasets, prompts, and hyperparameters used in our experiments are fully documented in Section 4.1 and Appendix C. These descriptions should allow researchers to replicate our experimental setup andresults without requiring additional resources beyond those specifed.
Appendix A Principle of Algorithm for Invert4TVG
Algorithm 1 GRPO Training with Randomized Invert-TVG Task Selection
1:Input: Video $V$, query $q$, LVLM parameters $\theta$, probability $p$, reward function $r_{\text{TVG}}(o)$, a list of Invert-TVG tasks $T_{\text{Invert-TVG}}=\{\text{Invert-TVG}_{i}\}_{i=1}^{N}$ with corresponding rewards $\{r_{i}\}_{i=1}^{N}$, learning rate $\eta$.
2:Output: Optimized LVLM parameters $\theta$ balancing localization accuracy and video-language alignment.
3: Define forward TVG task: $\text{TVG}(V,q)\rightarrow\tau$
4: Define a list of Invert-TVG tasks, where each $\text{Invert-TVG}_{i}(V,\tau)\rightarrow q^{\prime}_{i}$
5:while not converged do
6:  Sample a random value $u\sim\text{Uniform}(0,1)$
7:  if $u\leq p$ then
8:   Select reward $r=r_{\text{TVG}}(o)$ {Optimize parameters related to $\tau$ (localization accuracy)}
9:   Compute gradient $\nabla_{\theta}r$ w.r.t. parameters affecting $\tau$
10:  else
11:   Randomly sample an Invert-TVG task $\text{Invert-TVG}_{i}\sim T_{\text{Invert-TVG}}$ {Select a task from the list}
12:   Generate its output $q^{\prime}_{i}\leftarrow\text{Invert-TVG}_{i}(V,\tau)$
13:   Select the corresponding reward $r=r_{i}(o)$ {Optimize for the sampled Invert-TVG task}
14:   Compute gradient $\nabla_{\theta}r$ w.r.t. parameters affecting $q^{\prime}_{i}$
15:  end if
16:  Update parameters: $\theta\leftarrow\theta+\eta\nabla_{\theta}r$ {Gradient ascent to maximize reward}
17:end while
18:Result: The model retains both localization accuracy (via $\tau$) and diverse video-language alignment (via various $q^{\prime}_{i}$), which are complementary.
Appendix B Theoretical Justification for Multi-Task RL in Invert4TVG
To demonstrate the advantages of incorporating the Invert-TVG task into the RL framework, we provide a theoretical analysis showing that the multi-task approach improves semantic alignment and generalization compared to single-task TVG training. We follow the Pareto optimality framework in multi-task reinforcement learning, adapted to our setting where the joint reward balances temporal localization and semantic fidelity.
Let $\pi_{\theta}$ denote the policy (LVLM), and $D$ the data distribution over videos $V$, queries $q$, and ground-truth segments $\tau$. The single-task objective (TVG-only, as in prior works like Time-R1) maximizes:
$\max_{\pi_{\theta}}\mathbb{E}_{o\sim\pi_{\theta}}[R_{\text{TVG}}(o)]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}}),$ (14)
where $R_{\text{TVG}}(o)=r_{\text{IoU}}(o)+r_{\text{form}}(o)$.
In our multi-task setting, we introduce the joint reward $R_{\text{joint}}(o)=R_{\text{TVG}}(o)+\lambda R_{\text{Invert}}(o)$, with $\lambda>0$ balancing the tasks. The objective becomes:
$\max_{\pi_{\theta}}\mathbb{E}_{o\sim\pi_{\theta}}[R_{\text{joint}}(o)]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}}).$ (15)
Lemma 1 (Semantic Alignment Improvement).
The Invert-TVG task minimizes a semantic loss $L_{\text{sem}}=\mathbb{E}_{(V,\tau)\sim D}[d(q^{\prime},q)]$, where $d(\cdot,\cdot)$ is a distance metric (e.g., verb matching or KL divergence on embeddings). Then, the joint loss satisfies $L_{\text{joint}}\leq L_{\text{TVG}}+C$ for some constant $C>0$, as $R_{\text{Invert}}$ provides positive feedback on semantic fidelity.
Proof.
By Jensen’s inequality and non-negativity of $R_{\text{Invert}}\geq 0$ (binary rewards in our design), $\mathbb{E}[R_{\text{joint}}]\geq\mathbb{E}[R_{\text{TVG}}]+\lambda\min R_{\text{Invert}}\geq\mathbb{E}[R_{\text{TVG}}]$, assuming $R_{\text{Invert}}\geq 0$. This implies the multi-task policy reduces semantic drift, as Invert rewards enforce alignment (e.g., verb recovery). ∎
theorem 1 (Pareto Superiority).
The multi-task policy $\pi^{*}_{\text{joint}}$ is Pareto superior to the single-task policy $\pi^{*}_{\text{TVG}}$ if there exists $\theta$ such that $R_{\text{TVG}}(\pi_{\theta})\geq R_{\text{TVG}}(\pi^{*}_{\text{TVG}})$ and $R_{\text{Invert}}(\pi_{\theta})>0$.
Proof.
Consider the convex optimization formulation: minimize $L_{\text{TVG}}+\lambda L_{\text{Invert}}$. Assuming $L_{\text{Invert}}$ is convex (e.g., cross-entropy-like semantic loss), the Pareto frontier dominates the single-task optimum. The KL regularizer ensures the multi-task solution lies on a superior frontier, as feedback from Invert reduces divergence: $D_{\text{KL}}(\pi_{\text{joint}}\|\pi_{\text{ref}})\leq D_{\text{KL}}(\pi_{\text{TVG}}\|\pi_{\text{ref}})-\Delta$ for $\Delta>0$ from semantic regularization. ∎
Corollary 1 (Generalization Bound).
In migrating TVG to LVLMs, semantic drift is reduced by Invert tasks, yielding a generalization error bound: $\text{Err}_{\text{joint}}\leq\text{Err}_{\text{TVG}}-\eta\lambda$, where $\eta$ is a learning rate factor derived from multi-task boosting.
This analysis justifies the inclusion of Invert-TVG, showing improved alignment and generalization on the Pareto front.
Appendix C More Implementation details
We implement our model using the Qwen2.5-VL model as the backbone, selected for its robust feature extraction capabilities in video understanding tasks. To balance training efficiency and memory constraints, we sample video frames at 2 frames per second (FPS), adaptively resizing each frame to maintain approximately 2.8 million pixels per video. For example, a 50-second video yields 100 frames, each with a resolution of approximately 96 × 96 × 3 pixels. During the reinforcement fine-tuning phase, we train the model for 2 epochs with a batch size of 4. All experiments are conducted on a cluster equipped with eight NVIDIA A100 GPUs (40GB memory each), using CUDA 11.8 and Python 3.10. For natural language processing tasks, we employ the en_core_web_sm-3.8.0 model from the SpaCy library (12MB) to extract verbs from sentences. Random numbers between 0 and 1 are generated using numpy.random. The model is optimized using the AdamW optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.0. The learning rate is set to $5\times 10^{-5}$ . Training requires about 80 hours. All code, configurations, and preprocessing scripts are provided in the supplementary materials to ensure reproducibility.
Below, we provide a detailed breakdown of the computational overhead introduced by the three auxiliary tasks (data processing, reward computation, and extra forward/backward passes), including concise cost tables.
TVG VD VC AR Total
65h 6.5h 5h 3.5h 80h
As can be seen from the table, in our method, TVG overall consumed 65 hours (accounting for 81.25%), while Invert-TVG overall consumed 15 hours (accounting for 18.75%), with VD, VC, and AR consuming 6.5 hours, 5 hours, and 3.5 hours, respectively. It is important to note that if the Invert-TVG tasks were not used and only TVG tasks were employed, the total training time would still amount to 80 hours. The Invert-TVG task does not lead to an increase in training time, and our method just took the training time originally meant for the TVG task and used it for the Invert-TVG task.
Detailed time usage in the three auxiliary tasks per training step is shown in the following table:
Data Processing Forward Reward Backward
VC Offline* 117ms 81ms 89ms
VD Offline 121ms 46ms 90ms
AR Offline 119ms 10ms 90ms
*The data processing was fulfilled before the training.
Similarly, the VRAM costs are nearly identical to the baseline, with no additional overhead.
Model VRAM costs
Qwen-vl-2.5-3B 146G
Our 142G
Appendix D Ablation study
Figure 6: After Invert-SFT, the mIoU of Qwen2.5-VL-3B model and our Invert4TVG method are compared with those without Invert SFT.
Impact of Invert-SFT on Model Training. Invert-SFT refers to feeding ground-truth video clips into the model and requiring the model to output the corresponding event description based on these clips. The ground-truth clips are directly cropped from annotated temporal segments in temporal video grounding datasets, while the ”corresponding event” is the query to be localized. After Invert-SFT, the model’s outputs become more stable, facilitating subsequent training for Invert tasks. As shown in Figure 6, for the Qwen2.5-VL-3B model initialized with Invert-SFT, the mean Intersection-over-Union (mIoU) slightly decreased from 18.0 to 17.8 initially. However, after sufficient training, both Invert4TVG models, with and without Invert-SFT, reached convergence, achieving identical mIoU scores of 57.9.
Appendix E Qualitative Result
Figure 7: success case 1
Time localization of similar actions. As shown in Figure 7, our method can more accurately identify similar actions, such as opening and closing doors. Many other models have insufficient understanding of similar actions, such as picking up and putting things in a box. In the model’s judgment, it is likely to be classified as the same action because the model does not fully recognize the state in which the action occurs and continues. Our method can help the model understand actions more deeply and distinguish similar actions.
Figure 8: success case 2
Time localization of complex actions. As shown in Figure 8, our model recognizes the completion of the door closing action better. For some complex events, the time point we need to locate may occur at the completion of the action, rather than the beginning of the action. Other models sometimes consider the beginning of the action as the time point to be located, and then proceed with subsequent positioning from this time point. Our model has a clearer understanding of the start and end of the action and can effectively locate the time period when the action is in such a state of completion. Meanwhile, a correct understanding of the beginning and end of an action is also helpful for contextual reasoning.
Figure 9: success case 3
Situations involving multi-person. As shown in Figure 9, our method demonstrates a strong capability in handling situations involving multiple persons. When two individuals appear in the frame, other models are often susceptible to interference from the secondary person, leading to extended temporal localization periods. In contrast, our model possesses a more accurate understanding of actions, enabling it to precisely identify the core action to be localized and the state of the target person, thereby achieving superior results.
Figure 10: success case 4
Situations involving non-human object. As shown in Figure 10, our method is capable of handling scenarios where no person is present in the frame, such as when only a bonfire is shown and an eggplant is thrown into the fire. Other models, upon recognizing the keyword ”fire,” tend to predict very short temporal segments. Even if these predictions are accurate in timing, their Intersection over Union (IoU) remains low. In contrast, our approach focuses on understanding the action itself, resulting in predicted segments that are longer and closer to the ground truth temporal annotations.
Figure 11: success case 5
A person performing multiple actions. As shown in Figure 11, the event to be localized involves a person reading a book while standing up. Other models focus only on a single action, namely ”reading,” whereas our model disambiguates the actions, accurately identifying both the reading and the simultaneous act of standing up.
Figure 12: success case 6
A person performing multiple actions. As shown in Figure 12, the query contains a temporal cue such as “and then.” While other models treat the consecutive actions as a single event and attend more to the earlier action, our model recognizes both actions and their temporal order, yielding a more accurate localization.
Figure 13: success case 7
A person performing multiple actions. As shown in Figure 13, the event to be localized involves a causal relationship: a person in the video first sneezes and then takes medicine. Other models fail to accurately recognize the action of sneezing, leading them to rely on speculation and only localize the action of taking medicine. In contrast, our model successfully identifies both sneezing and taking medicine, understands the causal relationship between them, and achieves superior temporal localization results.
Figure 14: failure case
Deep understanding of event semantics. As shown in Figure 14, our method is not sensitive enough to some qualifiers, such as ”again” representing the second occurrence of an action, which requires the model to accurately identify the action while also accurately finding the time period during which the second action occurred. Our method, as well as other models, has some shortcomings in this aspect. When locating time, we may find the first occurrence time as the final answer. This is because the model does not have a deep understanding of the meaning of qualifiers in the query and fully considers it when locating video time.
Appendix F USE OF LARGE LANGUAGE MODELS
We used large language models (LLMs) in a limited and auxiliary manner during the preparation of this paper. Specifically, LLMs were employed to improve the fluency and readability of the manuscript by polishing grammar and style, without altering the technical content. Importantly, LLMs were not involved in formulating research ideas, designing methods, conducting experiments, analyzing results, or drawing conclusions. All technical contributions of this paper are solely the work of the authors.